A boy lives at X and wants to go to school at Z. From his home to X he has to first reach Y



and then Y to Z. He may go X to Y by either 3 bus routes or 2 train routes. From there, he can



either choose 4 bus routes or 5 train routes to reach Z. How many ways are there to go from X to



Z?

1
Expert's answer
2022-04-21T10:15:28-0400

There are 3+2=5 to reach Y from X. There are 4+5=9 ways to reach Z from Y. So, there are 5*9=45 ways to reach Z from X